verify the identity
cos^2 B-sin^2 B=1-2sin^2 B

Found 2 solutions by lwsshak3, w_parminder:
Answer by lwsshak3(11628)

verify the identity
cos^2 B-sin^2 B=1-2sin^2 B
start with left side:
cos^2 B-sin^2 B
1-sin^2B-sin^2B=1-2sin^2B
verified: left side=right side

Answer by w_parminder(53)

cos^2 B-sin^2 B=1-2sin^2 B
= (1-Sin^2 B)-Sin^2 B {Because Sin^2 B+Cos^2 B=1}
= 1-Sin^2 B - Sin^2 B
= 1-2Sin^2 B